[Brahms-l] Shift report 20050126 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Jan 26 2005 - 16:03:50 EST
1. Had 2 stores during this shift: from 8 to 10:30 (from night shift), and the second from 13:30 until 16:00 (still on, estimated time for dump, 18:30). 

2. Data runs: 

12505 - FS 8@B1219 (90k - tr. 2), MRS 90@B2255 (56k - tr. 3); 
12506 - FS 8@B608 (151k - tr. 2), MRS 90@B2255 (52k - tr. 3); 
12510 - only MRS 90@B2255 (22k - tr. 3) (the magnet people worked on D3); 
12511 - FS 8@B608 (100k - tr. 2), MRS 90@B2255 (56k - tr. 3); 
12512 - FS 8@B1723 (52k - tr. 2), MRS 90@B2255 (46k - tr. 3). 

Pedestal runs: 12507 (non-TPC pedestal), 12508 (TPC pedestal run). 

3. No detector problems found. Dana fixed the TOFW2 slat 38. 

Finished schedule for Fs8 for B polarities.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Jan 26 16:04:15 2005

This archive was generated by hypermail 2.1.8 : Wed Jan 26 2005 - 16:04:39 EST